# A Pairwise Comparison Relation-assisted Multi-objective Evolutionary Neural Architecture Search Method with Multi-population Mechanism

### v2

@@ IV-B Results on Standard Datasets
In this section, we show the results of our method on benchmark datasets to validate its effectiveness. We conducted the search and trained the final architectures separately on three datasets, which is different from many existing NAS methods. They mostly adopt a transfer learning setup where architectures searched on one dataset are directly applied to others. We think the transferred architectures may not be optimal on another dataset. Then, in order to validate the effectiveness of the proposed method, we compare the non-dominated architectures obtained by SMEM-NAS with those obtained by other state-of-the-art NAS methods. The selected peer methods can be broadly divided into three categories: manually designed by human experts, EA-based, and non-EA-based (e.g., RL and GD). In addition, we give the searched results by SMEM-NAS without the multi-population mechanism.
@@
Results on CIFAR-10: After completing the search using our method, we selected three promising architectures, and for ease of representation we named them SMEMNAS-S/M/L (sorted by MAdds). Fig. 5 and Table I show the results of SMEM-NAS and other methods on CIFAR-10. Fig. 5 visualizes the dominant relationship between the network models in terms of accuracy and model complexity. It is obvious that our models always outperform most other models. Table I shows more details. Firstly, from the perspective of time, our method takes 1.25 GPU days and significantly reduces the search time compared with methods based on EA and RL. The surrogate model we used effectively accelerates the search process. Compared with the GD-based method, although there are some deficiencies in time, the architecture we searched has better performance. Our model achieves the accuracy of 98.13%, surpassing almost all peer competitors. Among them, our model is slightly lower than FairNAS-A, but has nearly 100M lower MAdds and the search time is much less than FairNAS. In general, SMEM-NAS is superior to or consistent with other models in different metrics. Moreover, we also present the results obtained without employing the multipopulation mechanism (denoted as w/o), thereby demonstrating the effectiveness of this mechanism.
Figure 6: Accuracy and number of multi-adds in millions on CIFAR-100. Models from multi-objective approaches are joined with lines.
Results on CIFAR-100: Similar to the settings on CIFAR-10, we also conducted experiments on CIFAR-100. Fig. 6 and Table II show the results of our model and other methods. We also provide three models on this dataset, i.e., large, medium and small, in which the highest accuracy is 87.95%. From Fig. 6, it is clearly observed that our architectures dominate most methods in terms of accuracy and complexity. Table II shows more details. On this dataset, although our results are slightly inferior to EfficientNet, the hand-crafted model requires significant expertise and labour. Additionally, we can see that the architecture with the highest accuracy searched on the CIFAR-10 dataset transferred to CIFAR-100 does not perform optimally, achieving 87.18% accuracy. Moreover, the search time of our method on this dataset is shorter than that of all other methods that also employ evolutionary algorithms.
TABLE III: Comparison with state-of-the-art image classifiers on the ImageNet dataset. The search cost excludes the supernet training cost.
@@
EfficientNet-B0 [13] 76.3 93.2 390 5.3 - - 2019 manual
PA-DARTS [57] 75.3 92.25 - 5.2 0.4 RTX 3090 2024 GD
RelativeNAS [60] 75.12 92.3 563 5.05 0.4 GTX 1080Ti 2023 GD
$\beta$-DARTS [72] 76.1 93.0 609 5.5 0.4 - 2022 GD
NAP [73] 75.5 92.6 574 4.8 4 RTX 2080Ti 2022 GD
iDARTS [54] 75.3 92.3 568 5.1 1.9 Tesla V100 2023 GD
SQNAS [74] 73.8 91.4 540 - 0.25 Tesla V100 2024 GD
Shapely-NAS [75] 76.1 - 582 5.4 4.2 - 2022 GD
RandomNAS [76] 74.2 - - 5.9 0.5 - 2025 GD
SPNAS [66] 78.62 94.07 687 6.6 0.37 RTX 3090 2025 EA
OnceForAll [51] 76.9 93.2 230 - 2 Tesla V100 2020 EA
FairNAS [49] 77.5 - 392 5.9 12 - 2021 EA
EEEA-Net-C2 [63] 76.2 92.7 217 6.0 - - 2021 EA
NSGANetV2 [25] 77.5 93.5 225 6.1 1 - 2020 EA
MixPath [77] 77.2 93.5 378 5.1 10.3 Tesla V100 2023 EA
DPS [78] 75.6 92.7 578 5.3 0.06 RTX A6000 2023 EA
CENAS-C [79] 79.6 94.4 482 4.98 1.98 RTX 2080Ti 2024 EA
SLE-NAS-B [80] 75.7 92.5 412 4.5 3 RTX A6000 2023 EA
MnasNet [20] 76.7 93.3 403 5.2 4.5 64*TPUv2 2019 RL
BNAS [69] 74.3 92.2 - 3.9 0.19 - 2022 RL
DBNAS [70] 77.6 93.5 386 4.9 0.9 RTX 3090 2025 RL
@@
Results on ImageNet: Different from the conclusion on the CIFAR datasets, there is no need for pre-training before searching on the ImageNet dataset. Table III shows the results of the proposed method and other comparative methods. SMEM-NAS outperforms most other methods and achieves the accuracy of 78.91% with the MAdds of 570M. In addition, our proposed surrogate model effectively reduces the search time, and only takes 4 hours to complete the search. It is worth mentioning that although the final accuracy achieved by our method is somewhat lower than that of CENAS-C, our search time and resource consumption are significantly lower than CENAS-C. We obtained the final results using only 10% of the search time with just one GPU.
In summary, the results on the three benchmark datasets validate the effectiveness of the proposed method in terms of the architectural performance and time consumption.
@@
Figure 7: The comparison of results between different probabilities for using two populations on the ImageNet. Different probability values indicate the probability that the parents generating the next generation of offspring will come from the two populations. We plot the distribution of architectures in the archive during the search process in three figures. The horizontal and vertical axes are MAdds and error, respectively.
@@
Figure 8: The comparison between SMEM-NAS with and without the multi-population mechanism (MP) on CIFAR-100. We plot the distribution of architectures in the archive during the search process in two figures. The blue dots represent all the candidate architectures searched in the 5th iteration, and the red dots represent the new candidates generated till the 20th iteration. In addition, the blue and red lines represent the pareto front of the 5th iteration and 20th iteration respectively.
Figure 9: ParetoFrontUpdate: Changes of the pareto front during the search process on CIFAR-100.
(a) SVM ($KTau$ = 0.93)
(b) MLP ($KTau$ = 0.91)
(c) KNN ($KTau$ = 0.88)
(d) RF ($KTau$ = 0.85)
Figure 10: Search with different surrogate models. The top row shows the results of the search using four different models. The horizontal and vertical axes are MAdds and error, respectively. The bottom row shows the correlation between the predicted accuracy ranking and the true ranking using $KTau$.
@@ IV-C Performance of Multi-population Mechanism
The multi-population mechanism is an important strategy to increase the diversity of the solutions while ensuring convergence in this study. In order to verify its effectiveness, more experiments were conducted on the CIFAR-100 dataset in this section. We separately conducted the architecture search with and without the multi-population mechanism, and the comparison between the two search processes is shown in Fig. 8. The figures show the changes of the pareto front and the distribution of new individuals generated during the evolutionary process. Fig. 8a shows the search process without the multi-population mechanism and Fig. 8b shows the search process using the multi-population mechanism. Fig. 8 shows the distributions of the architectures searched in the archive, which have been realistically evaluated. Specifically, in both figures, the blue dots represent individuals at the initial and early stages of the evolutionary process, and the corresponding blue line is the pareto front at the early stage. In Fig. 8a, it is obvious that except for the randomly initialized individuals, the new individuals generated in the later evolution are all almost distributed in the same solution region, and there is no obvious change in the pareto front. However, in Fig. 8b, we can see that the new individuals generated are uniformly distributed unlike without the multi-population mechanism. The pareto front is also further updated in the excellent direction (lower error rate, smaller model size) and more promising individuals are explored. Evidently, the multi-population mechanism is better able to ensure the diversity of solutions and prevent them from falling into the local optima. Additionally, we show changes of the pareto front during the search process by using our method in Fig. 9. As we can see that the pareto front has stabilized in the early stage of the search, and is constantly changing and updating with the increase of iterations. In general, that is the same as our expected effect, accelerating the convergence speed while maintaining the diversity.
IV-D Ablation Study
@@
Comparison of classification models: In order to construct more accurate surrogate models, we considered four classification models: SVM, random forest, k-nearest neighbors, and multi-layer perceptron. We performed four groups of experiments on the CIFAR-100 dataset. For more intuitive comparisons, we computed the correlation of 300 fully trained and evaluated architectures during the search process with each surrogate model. We choose the Kendall’s Tau coefficient ($KTau$) to show the correlation between their true rankings and predicted rankings. Fig. 10 shows the searched results using different models and the corresponding $KTau$. Architectures from the archive were used to train surrogate models. After that, new $K$ architectures were retained in each generation and used for testing. The figure shows the results of the last generation. From the $KTau$ and the distribution of individuals in the search process, it can be seen that SVM is the best classifier with very limited training samples. The new individuals generated during the evolution when using KNN and RF as surrogate models are worse and non-uniformly distributed, which may be caused by the inaccurate prediction at the early stage. In addition, the performance of MLP is similar to SVM as shown in the figure. But, SVM can take less time for training and achieve better results than MLP with the limited samples. Therefore, we choose SVM as the surrogate model.
@@
Compared to traditional surrogate models, we conducted data augmentation, expanding from $n$ samples to $n(n-1)/2$, which greatly increases the training data. Increased training data can effectively improve the reliability of the surrogate model. Table IV shows the performance comparison of the proposed surrogate model with the traditional regression surrogate models. We used 300 samples for training models, and in each experiment each surrogate model was trained with the same data. We trained three models with the proposed classification-based method (SVM, RF, and KNN), and we trained another three models with the traditional regression-based method (MLP, Decision Tree, and AdaBoost). Table IV highlights that while the training and prediction time of the proposed surrogate model surpasses that of traditional models, classification-based surrogate models demonstrate superior performance in terms of correlation coefficients. Among these, our SVM achieves the highest $KTau$ coefficient at 0.8113. Therefore, the proposed surrogate model can provide more reliable prediction results and is more capable of retaining and choosing the excellent individuals during the search. Moreover, although it takes more time to train the proposed surrogate model, the training and prediction only take a few more minutes during the actual search, which is acceptable compared to the whole search time.
TABLE IV: Comparison with different surrogate models. We record the Spearman coefficients ($\rho$), Kendall’s Tau coefficients ($KTau$) between the predicted accuracy rankings and the true accuracy rankings. The results are the average of ten experiments carried out.
@@
TABLE V: The hypervolume values are achieved with different size of initial samples.
HV Initial size
0.8313 25
0.8339 50
0.8363 100
0.8360 150
TABLE VI: Comparison with different sizes of initial samples to train surrogate model.
$KTau$ $\rho$ Initial size
0.4465 0.6154 25
0.5846 0.7485 50
0.6955 0.8575 100
0.7472 0.9123 150
Hyper-parameters: The initial population is important for both the training of the surrogate model and the subsequent evolution. In order to achieve a better balance between search efficiency and search quality, we tested the appropriate value of the initial population size, $N$, by conducting the following small-scale experiments on the CIFAR-10. We set different sizes of the initial population, i.e., $N=25$, $50$, $100$ and $150$. We recorded the effect of different population sizes on the hypervolume (HV) value and the comparison results are shown in Table. VI. From the table, it can be seen that HV achieves the highest value when the population size $N=100$. Although the HV obtained at $N=150$ and at $N=100$ are almost equivalent, 50% more individuals need to be evaluated when $N=150$, which is more time consuming. In addition, we tested the prediction performance of the initial surrogate model as well. As seen in Table VI, the reliability of the surrogate model is low if the initial population size is too small, which is another reason for the unfavourable search results at $N=25$ and $50$. Besides, if the surrogate model has low reliability, the search results are prone to be less stable. Therefore, after comprehensive consideration, we choose $N=100$ as the size of the initial population.
@@ V-A Conclusion
In this paper, we propose an evolutionary multi-objective algorithm by using a multi-population mechanism and a surrogate model based on pairwise comparison relations to perform neural architecture search. The proposed method utilizes a surrogate model to predict accuracy rankings of architectures without focusing on accuracy values. The proposed surrogate model makes full use of the limited training samples and can greatly reduce the search overhead. Besides, a multi-population mechanism is used to improve the diversity in multi-objective NAS, which not only increases the population diversity in the evolution process, but also accelerates the convergence of the algorithm. We have conducted a series of experiments on benchmark datasets to verify the effectiveness of SMEM-NAS. The final architectures found by SMEM-NAS outperform most existing search methods.
V-B Limitations and Future Work
@@
In this work, our surrogate model only considers the accuracy and ignores other indicators of networks. Single-objective surrogate models are not effectively adapted to multi-objective search frameworks. Our next work plans to explore the multi-objective surrogate model to predict domination relations of architectures. Besides, it can be observed that the initial population is vital to the results of evolutionary algorithms. The proposed method simply uses random sampling to initialise the population, and the sampled architectures are non-uniformly distributed in the objective space, which may not be a proper approach for subsequent evolutionary exploration. Therefore, we plan to investigate efficient sampling methods to initialize the population in future work by considering a given problem’s features. In addition, to test the performance of the model and its applicability in different computer vision tasks, we will also use a pretrained model on ImageNet as the backbone to test its performance on other tasks such as object detection and semantic segmentation.

### v3

@@ IV-B Results on Standard Datasets
In this section, we show the results of our method on benchmark datasets to validate its effectiveness. We conducted the search and trained the final architectures separately on three datasets. Then, in order to validate the effectiveness of the proposed method, we compare the non-dominated architectures obtained by SMEMNAS with those obtained by other state-of-the-art NAS methods. The selected peer methods can be broadly divided into three categories: manually designed by human experts, EA-based, and non-EA-based (e.g., Bayesian, RL and GD). In addition, we give the searched results by SMEMNAS without the multi-population mechanism. The results on CIFAR-10 and ImageNet are presented in Tables I and II, and the results in CIFAR-100 are presented in the supplementary materials due to the page limitation.
@@
Results on CIFAR-10: After completing the search using our method, we selected three promising architectures, and for ease of representation we named them SMEMNAS-S/M/L (sorted by MAdds). Fig. 5 and Table I show the results of SMEMNAS and other methods on CIFAR-10. Fig. 5 visualizes the dominant relationship between the network models in terms of accuracy and model complexity. It is obvious that our models always outperform most other models. Table I shows more details. Firstly, from the perspective of time, our method takes 1.25 GPU days and significantly reduces the search time compared with methods based on EA and RL. The surrogate model we used effectively accelerates the search process. Compared with the GD-based method, although there are some deficiencies in time, the architecture we searched has better performance. Our model achieves the accuracy of 98.13%, surpassing almost all peer competitors. Among them, our model is slightly lower than FairNAS-A, but has nearly 100M lower MAdds and the search time is much less than FairNAS. In general, SMEMNAS is superior to or consistent with other models in different metrics. Moreover, we also present the results obtained without employing the multipopulation mechanism (denoted as w/o), thereby demonstrating the effectiveness of this mechanism.
TABLE II: Comparison with state-of-the-art image classifiers on the ImageNet dataset. The search cost excludes the supernet training cost.
@@
EfficientNet-B0 [tan_efficientnet_2019] 76.3 93.2 390 5.3 - - 2019 manual
Optuna (so) [akiba_optuna_2019] 78.61 93.98 732 7.66 0.19 RTX 3090 2019 Bayesian
Optuna-M (mo) [akiba_optuna_2019] 77.48 93.3 329 5.29 0.16 RTX 3090 2019 Bayesian
Optuna-L (mo) [akiba_optuna_2019] 78.63 94.01 675 7.66 0.16 RTX 3090 2019 Bayesian
PA-DARTS [xue_improved_2024] 75.3 92.25 - 5.2 0.4 RTX 3090 2024 GD
RelativeNAS [tan_relativenas_2023] 75.12 92.3 563 5.05 0.4 GTX 1080Ti 2023 GD
$\beta$-DARTS [ye_-darts_2022] 76.1 93.0 609 5.5 0.4 - 2022 GD
NAP [ding_nap_2022] 75.5 92.6 574 4.8 4 RTX 2080Ti 2022 GD
iDARTS [wang_idarts_2023] 75.3 92.3 568 5.1 1.9 Tesla V100 2023 GD
SQNAS [gm_sequential_2024] 73.8 91.4 540 - 0.25 Tesla V100 2024 GD
Shapely-NAS [xiao_shapley-nas_2022] 76.1 - 582 5.4 4.2 - 2022 GD
RandomNAS [ma_defying_2025] 74.2 - - 5.9 0.5 - 2025 GD
SPNAS [jiang_score_2025] 78.62 94.07 687 6.6 0.37 RTX 3090 2025 EA
Once-For-All [cai_once_2020] 76.9 93.2 230 - 2 Tesla V100 2020 EA
FairNAS [chu_fairnas_2021] 77.5 - 392 5.9 12 - 2021 EA
NSGANetV2 [lu_nsganetv2_2020] 77.5 93.5 225 6.1 1 - 2020 EA
MixPath [chu_mixpath_2023] 77.2 93.5 378 5.1 10.3 Tesla V100 2023 EA
DPS [li_extensible_2023] 75.6 92.7 578 5.3 0.06 RTX A6000 2023 EA
CENAS-C [ma_pareto-wise_2024] 79.6 94.4 482 4.98 1.98 RTX 2080Ti 2024 EA
SLE-NAS-B [huang_split-level_2023] 75.7 92.5 412 4.5 3 RTX A6000 2023 EA
MnasNet [tan_mnasnet_2019] 76.7 93.3 403 5.2 4.5 64*TPUv2 2019 RL
BNAS [ding_bnas_2022] 74.3 92.2 - 3.9 0.19 - 2022 RL
DBNAS [yang_deeply_2025] 77.6 93.5 386 4.9 0.9 RTX 3090 2025 RL
@@
Results on ImageNet: Different from the conclusion on the CIFAR datasets, there is no need for pre-training before searching on the ImageNet dataset. Table II shows the results of the proposed method and other comparative methods. SMEMNAS outperforms most other methods and achieves the accuracy of 78.91% with the MAdds of 570M. In addition, our proposed surrogate model effectively reduces the search time, and only takes 4 hours to complete the search. It is worth mentioning that although the final accuracy achieved by our method is somewhat lower than that of CENAS-C, our search time and resource consumption are significantly lower than CENAS-C. We obtained the final results using only 10% of the search time with just one GPU. Additionally, we conducted experiments using the Optuna method with the same number of true evaluations on our search space. In Table II, we present the final training results, and it can be observed that even within the same search space, the highest-accuracy architecture obtained by SMEMNAS surpasses the classification accuracy achieved by the Optuna method while reducing MAdds by 100M. For a more detailed comparison between SMEMNAS and Optuna methods, please refer to the supplementary materials.
@@
Figure 6: The comparison of results between different probabilities for using two populations on the ImageNet. Different probability values in the captions of sub-figures indicate the probability that the parents generating the next generation of offspring will come from the two populations. We plot the distribution of architectures in the archive during the search process in three figures. The horizontal and vertical axes are MAdds and error, respectively.
@@
Figure 7: The comparison between SMEMNAS with and without the multi-population mechanism (MP) on CIFAR-100. We plot the distribution of architectures in the archive during the search process in two figures.
(a) SVM (KTauKTau = 0.93)
(b) MLP (KTauKTau = 0.91)
(c) KNN (KTauKTau = 0.88)
(d) RF (KTauKTau = 0.85)
Figure 8: Search with different surrogate models. The top row shows the results of the search using four different models. The horizontal and vertical axes are MAdds and error, respectively. The bottom row shows the correlation between the predicted accuracy ranking and the true ranking using KTauKTau.
@@ IV-C Performance of Multi-population Mechanism
The multi-population mechanism is an important strategy to increase the diversity of the solutions while ensuring convergence in this study. In order to verify its effectiveness, more experiments were conducted on the CIFAR-100 dataset in the supplementary materials. We separately conducted the architecture search with and without the multi-population mechanism, and the comparison between the two search processes is shown in Fig. 7. The figures show the distributions of the architectures searched in the archive and new individuals generated during the evolutionary process. Fig. 7a shows the search process without the multi-population mechanism and Fig. 7b shows the search process using the multi-population mechanism. Specifically, in both figures, the blue dots represent all the candidate architectures searched in the 5th iteration, and the red dots represent the new candidates generated till the 20th iteration. In addition, the blue and red lines represent the Pareto front of the 5th iteration and 20th iteration respectively. In Fig. 7a, it is obvious that except for the randomly initialized individuals, the new individuals generated in the later evolution are all almost distributed in the same solution region, and there is no obvious change in the Pareto front. However, in Fig. 7b, we can see that the new individuals generated are uniformly distributed unlike without the multi-population mechanism. The Pareto front is also further updated in the excellent direction (lower error rate, smaller model size) and more promising individuals are explored. Evidently, the multi-population mechanism is better able to ensure the diversity of solutions and prevent them from falling into the local optima.
IV-D Ablation Studies
@@
The surrogate model is one of the critical components of SMEMNAS. Besides, the hyper-parameters in multi-population mechanism also affect the performance. We now present ablation studies examining different design choices for the surrogate model and other key hyper-parameters. Comparison of classification models: In order to construct more accurate surrogate models, we considered four classification models: SVM, random forest, k-nearest neighbors, and multi-layer perceptron. We performed four groups of experiments on the CIFAR-100 dataset. For more intuitive comparisons, we computed the correlation of 300 fully trained and evaluated architectures during the search process with each surrogate model. We choose the Kendall’s Tau coefficient ($KTau$) to show the correlation between their true rankings and predicted rankings. Fig. 8 shows the searched results using different models and the corresponding $KTau$. Architectures from the archive were used to train surrogate models. After that, new $K$ architectures were retained in each generation and used for testing. The figure shows the results of the last generation. From the $KTau$ and the distribution of individuals in the search process, it can be seen that SVM is the best classifier with very limited training samples. The new individuals generated during the evolution when using KNN and RF as surrogate models are worse and non-uniformly distributed, which may be caused by the inaccurate prediction at the early stage. In addition, the performance of MLP is similar to SVM as shown in the figure. But, SVM can take less time for training and achieve better results than MLP with the limited samples. Therefore, we choose SVM as the surrogate model.
@@
Compared to traditional surrogate models, we conducted data augmentation, expanding from $n$ samples to n(n1)/2n(n-1)/2, which greatly increases the training data. Increased training data can effectively improve the reliability of the surrogate model. Table III shows the performance comparison of the proposed surrogate model with the traditional regression surrogate models. We used 300 samples for training models, and in each experiment each surrogate model was trained with the same data. We trained three models with the proposed classification-based method (SVM, RF, and KNN), and we trained another three models with the traditional regression-based method (MLP, Decision Tree, and AdaBoost). Table III highlights that while the training and prediction time of the proposed surrogate model surpasses that of traditional models, classification-based surrogate models demonstrate superior performance in terms of correlation coefficients. Among these, our SVM achieves the highest $KTau$ coefficient at 0.8113. Therefore, the proposed surrogate model can provide more reliable prediction results and is more capable of retaining and choosing the excellent individuals during the search. Moreover, although it takes more time to train the proposed surrogate model, the training and prediction only take a few more minutes during the actual search, which is acceptable compared to the whole search time.
TABLE III: Comparison with different surrogate models. We record the Spearman coefficients ($\rho$), Kendall’s Tau coefficients (KTauKTau) between the predicted accuracy rankings and the true accuracy rankings. The results are the average of ten experiments carried out.
@@
TABLE IV: The hypervolume values are achieved with different size of initial samples.
Initial size 25 50 100 150
HV 0.8313 0.8339 0.8363 0.8360
TABLE V: Comparison with different sizes of initial samples to train surrogate model.
Initial size 25 50 100 150
KTauKTau 0.4465 0.5846 0.6955 0.7472
$\rho$ 0.6154 0.7485 0.8575 0.9123
Initial population size: The initial population is important for both the training of the surrogate model and the subsequent evolution. In order to achieve a better balance between search efficiency and search quality, we tested the appropriate value of the initial population size, $N$, by conducting the following small-scale experiments on the CIFAR-10. We set different sizes of the initial population, i.e., $N=25$, 5050, 100100 and $150$. We recorded the effect of different population sizes on the hypervolume (HV) value and the comparison results are shown in Table. IV. From the table, it can be seen that HV achieves the highest value when the population size $N=100$. Although the HV obtained at $N=150$ and at $N=100$ are almost equivalent, 50% more individuals need to be evaluated when $N=150$, which is more time consuming. In addition, we tested the prediction performance of the initial surrogate model as well. As seen in Table V, the reliability of the surrogate model is low if the initial population size is too small, which is another reason for the unfavorable search results at $N=25$ and $50$. Besides, if the surrogate model has low reliability, the search results are prone to be less stable. Therefore, after comprehensive consideration, we choose $N=100$ as the size of the initial population.
(a) Architectures distribution of static threshold
(b) Architectures distribution of SMEMNAS
Figure 9: The comparison between static threshold and dynamic threshold on ImageNet. We plot the distribution of architectures in the final archive in two figures. The blue dots represent all the candidate architectures searched in the iterations, and the red line represent the Pareto front.
Dynamic threshold: In SMEMNAS, the $Threshold$ that controls population generation is dynamically regulated. As can be observed from Eq. (4) and Eq. (7), when dynamic control is employed, even during phases that tend to favor diversity expansion, convergence is occasionally considered, and vice versa. To validate the effectiveness of this configuration, we designed a set of experiments with static control. We set the $Threshold$ value to consistently equal 0.3 in the early stage and 0.7 in the later stage, meaning that convergence is always considered in the early stage while diversity is always considered in the later stage. Fig. 9 demonstrates the results of the search Pareto front using static threshold compare with SMEMNAS. We can observe that the Pareto front of this configuration provides less diverse candidate solutions compared to the dynamic setting. Further analysis of dynamic threshold can be found in the supplementary materials due to page limitations.
@@ V Conclusion and Future Work
In this paper, we propose an evolutionary multi-objective algorithm that leverages a multi-population mechanism and a surrogate model based on pairwise comparison relations for neural architecture search. The proposed method utilizes a surrogate model to predict accuracy rankings rather than absolute accuracy values, effectively exploiting limited training samples and substantially reducing search overhead. Furthermore, a multi-population mechanism improves diversity in multi-objective NAS, increasing population diversity during evolution process, and accelerating algorithm convergence. Experiments on benchmark datasets verify the effectiveness of SMEMNAS. The final architectures discovered by SMEMNAS outperform most existing search methods.
@@
In this work, our surrogate model only considers accuracy and ignores other network indicators. Single-objective surrogate models are not effectively adapted to multi-objective search frameworks. Future work will explore the multi-objective surrogate models to predict domination relations among architectures. Besides, the initial population significantly influences evolutionary algorithm performance. The proposed method simply uses random sampling to initialize the population, and the sampled architectures are non-uniformly distributed in the objective space, which may not provide an optimal foundation for subsequent evolutionary exploration. Therefore, future investigations will focus on efficient initialization methods, including diverse sampling strategies and Bayesian approaches. Additionally, to test model performance and applicability in different computer vision tasks, we will use a pretrained model on ImageNet as the backbone to test performance on other tasks such as object detection and semantic segmentation. Furthermore, an important direction for future research is to systematically investigate the transferability of architectures discovered by SMEMNAS across significantly different domains and datasets. While our method demonstrates strong performance on image classification benchmarks (CIFAR-10, CIFAR-100, and ImageNet), understanding how well these architectures generalize to other visual recognition tasks, different data distributions, and even non-vision domains would provide valuable insights into our approach’s robustness and generalizability. This cross-domain transferability analysis would help establish the practical applicability of SMEMNAS in diverse real-world scenarios.
[Uncaptioned image] Yu Xue (Senior Member, IEEE) received the Ph.D. degree from the School of Computer Science and Technology, Nanjing University of Aeronautics and Astronautics, Nanjing, China, in 2013. He was a Visiting Scholar with the School of Engineering and Computer Science, Victoria University of Wellington, Wellington, New Zealand, from August 2016 to August 2017. He was a Research Scholar with the Department of Computer Science and Engineering, Michigan State University, East Lansing, MI, USA, from October 2017 to November 2018. He is currently a Professor with the School of Software, Nanjing University of Information Science and Technology, Nanjing. His research interests include deep learning, evolutionary computation, machine learning, computer vision, and feature map selection.
[Uncaptioned image] Pengcheng Jiang (Graduate Student Member, IEEE) received the B.E. degree from Nanjing University of Information Science and Technology, China, in 2020. He is currently pursuing the Ph.D. degree with the School of Software in Nanjing University of Information Science and Technology, China. His current research interests include feature selection, evolutionary computation, neural architecture search, and model compression.
[Uncaptioned image] Chenchen Zhu received the B.E. degree from Nanjing University of Information Science and Technology, China, in 2022, where she is currently pursuing a master’s degree. Her research interests include deep learning, multi-objective optimization and neural architecture search.
[Uncaptioned image] MengChu Zhou (Fellow, IEEE) received his Ph. D. degree from Rensselaer Polytechnic Institute, Troy, NY in 1990 and then joined New Jersey Institute of Technology where he has been Distinguished Professor since 2013. His interests are in Petri nets, automation, robotics, big data, Internet of Things, cloud/edge computing, and AI. He has 1200+ publications including 17 books, 850+ journal papers (650+ in IEEE transactions), 31 patents and 32 book-chapters. He is Fellow of IFAC, AAAS, CAA and NAI.
[Uncaptioned image] Mohamed Wahib is a team leader of the “High Performance Artificial Intelligence Systems Research Team” at RIKEN Center for Computational Science (R-CCS), Kobe, Japan. Prior to that he worked as is a senior scientist at AIST/TokyoTech Open Innovation Laboratory, Tokyo, Japan. He received his Ph.D. in Computer Science in 2012 from Hokkaido University, Japan. His research interests revolve around the central topic of highperformance programming systems, in the context of HPC and AI. He is actively working on several projects including high-level frameworks for programming traditional scientific applications, as well as high-performance AI.
[Uncaptioned image] Moncef Gabbouj (Fellow, IEEE) was an Academy of Finland Professor. He is currently a Professor of information technology with the Department of Computing Sciences, Tampere University, Finland. His research interests include big data analytics, multimedia analysis, artificial intelligence, machine learning, pattern recognition, nonlinear signal processing, video processing, and coding. He is a fellow of Asia–Pacific Artificial Intelligence Association. He is a member of the Academia Europaea, the Finnish Academy of Science and Letters, and the Finnish Academy of Engineering Sciences. He is the Finland Site Director of the NSF IUCRC funded Center for Big Learning.
See pages - of material_final.pdf